OpenStudy (amistre64):

How would we go about determining the continued fraction expansion of: \(\sqrt{5}\)?

OpenStudy (anonymous):

1+1/(5+1/(5+1/(5+1/5+...)))

OpenStudy (anonymous):

I'm wrong.... :-P

OpenStudy (amistre64):

yeah, the result is (2;4,4,4,...)

OpenStudy (amistre64):

all irrational algebraic numbers have a periodic continued fraction expansion. Which leaves us with transcendentals having an infinite expansion

OpenStudy (amistre64):

this is a slow site today ...

OpenStudy (zarkon):

\[\lfloor\sqrt{5}\rfloor=2\] \[\left\lfloor\frac{1}{\sqrt{5}-2}\right\rfloor=4\] \[\left\lfloor\frac{1}{\frac{1}{\sqrt{5}-2}-4}\right\rfloor=4\] ...

OpenStudy (anonymous):

try newton's method.

OpenStudy (zarkon):

continue this way will give the [2,4,4,4...]

OpenStudy (anonymous):

amistre, i found this if it helps; it is a special case